OpenStudy (anonymous):

What is the next term in the following sequence: 3, -15, 75, -375…? What is the rule for the following sequence: 8, 4, 2, 1…?

OpenStudy (welshfella):

for the first one calculate second term / first term and third term / second term what do you notice?

OpenStudy (kash_thesmartguy):

1. 1875 2. Dividing by 2.

OpenStudy (anonymous):

for the first one i think it is an = a1 * (1/2)(n-1), a1 = 8 btw

OpenStudy (welshfella):

@Kash_TheSmartGuy please don't give direct answers - read the code of conduct

OpenStudy (kash_thesmartguy):

Sorry

OpenStudy (anonymous):

But the second ones answer isn't on the choices

OpenStudy (welshfella):

the first one is a geometric sequence where the next term is obtained by multiplying by a common ratio

OpenStudy (welshfella):

the second one is also geometric - do you know what the nest number is?

OpenStudy (anonymous):

you mean next

OpenStudy (anonymous):

and I think it should be .5

OpenStudy (welshfella):

yes not nest!!

OpenStudy (welshfella):

right each term is formed by multiplying by 0.5 - same thing as dividing by 2 as Kash said

OpenStudy (welshfella):

and if you do the divisions i mentioned in my first post you'll get the common ratio for the first sequence

OpenStudy (welshfella):

and to get the next term you multiply -375 by this common ratio

OpenStudy (anonymous):

an = a1 * (1/2)(n-1), a1 =8 I guess maybe